OpenStudy (anonymous):

\[\sqrt[45]{\pi} =?\]

OpenStudy (anonymous):

\[\sqrt[45]{\pi}\]

OpenStudy (anonymous):

you need an approximate decimal answer?

OpenStudy (anonymous):

sure or fraction would be better

OpenStudy (anonymous):

There is no fraction.. Pi is not rational.

OpenStudy (anonymous):

i need to plug it into cos(x^45)

OpenStudy (anonymous):

ok then decimal

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=pi^%281%2F45%29

OpenStudy (anonymous):

heres what im doing....\[indefinite \int\limits x ^{45}\cos(x^{45}) dx \] becomes \[-1/45 \int\limits \cos (u) du\] i need to convert these in to the limit of \[\int\limits_{{0}}^{\sqrt[45]{\pi}}\] and if x = \[\sqrt[45]{\pi} then (u) = what?\]

OpenStudy (anonymous):

ok thanks..but how do it do the above?^^